A mother is four times as old as his son. In 6 years, she will be three as old as he is. How old is the son now?
Answer by macston(5194)

.
M=mother's age; S=son's age
.
M=4S Use this to substitute for M.
.
M+6=3(S+6) Substitute for M from above.
4S+6=3S+18
S=12
.
ANSWER: The son is 12 years old.
.
CHECK:
M=4(S)=4(12)=48yrs
Mother is 48 years old.
M+6=3(S+6)
48+6=3(12+6)
54=3(18)
54=54